if the lateral surface area of a cube is 100 square inches what is the length of an edge?
Answer by checkley77(12844)

There are 6 square surfaces on a cube.
6x^2=100
x^2=100/6
x^2=16.6667
x=sqrt16.6667
x=4.0825 in. on each side ans.